OpenStudy (anonymous):

If I know the lengths of two sides of a right triangle, how do I find the third?

OpenStudy (anonymous):

pythagorean thereom a^2+b^2=c^2

OpenStudy (nincompoop):

no thanks or no thank you?

OpenStudy (nincompoop):

a little diagram that goes with the "pythagorean theorem"  if we were to look only for a b or c, we should get the square root of the side that we were looking for. \[c=\sqrt{a^2 + b^2}\] \[b = \sqrt{c^2-a^2}\] \[a = \sqrt{c^2 - b^2}\]